# A Theoretical Framework for Data Efficient Multi-Source Transfer Learning Based on Cramér-Rao Bound

### v2

@@ Appendix B Proofs
Combining with Lemma  3.1, we will know that the K-L measure is
@@
B.4 Proof of Proposition 4.5
Similar to (B.2), we can get
$\mathbb{E}\left[D\left(P_{X;\hat{{\underline{\theta}}}}\middle\|P_{X;{\underline{\theta}}_{0}}\right)\right]$
@@
[…]
So we can know that $\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-E_{\hat{{\underline{\theta}}}}\right)^{2}\right]$ has the bound
@@
Combining (B.4), (54) and (55), we can know that the K-L measure is
@@
[…]
So we can know that $\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-E_{\hat{{\underline{\theta}}}}\right)^{2}\right]$ has the bound
@@
$E_{\hat{{\underline{\theta}}}}=\frac{N_{0}{\underline{\theta}}_{0}+\sum\limits_{i=1}^{k}n_{i}{\underline{\theta}}_{i}}{N_{0}+\sum\limits_{i=1}^{K}n_{i}}=\frac{N_{0}{\underline{\theta}}_{0}+\sum\limits_{i=1}^{k}n_{i}{\underline{\theta}}_{i}}{N_{0}+s}+O\left(\frac{1}{N_{0}+s}\right)$ (60)
Combining (B.2), (59) and(60) , we can know that the K-L measure is
$\frac{1}{2}\frac{1}{N_{0}+s}+\frac{1}{2}\frac{s^{2}}{(N_{0}+s)^{2}}J({\underline{\theta}}_{0})\left(\sum\limits_{i=1}^{k}\alpha_{i}({\underline{\theta}}_{i}-{\underline{\theta}}_{0})\right)^{2}+o\left(\frac{1}{N_{0}+s}\right)$ (61)
For the d-demension parameter, we can know that the K-L measure is $\underline{\theta}_{i}$
@@
Appendix C Experiment Details
C.1 Details information of LoRA framework experiments.
@@ Appendix C Experiment Details
C.2 Experimental Design and Model Adaptation
@@
C.3 Single-Source transfer performance details.
@@
C.4 Baselines experiments settings in Table 2.
Since the significant difference from unsupervised methods and supervised methods, the results of MSFDA(Shen et al., 2023), DATE(Han et al., 2023) and M3SDA(Peng et al., 2019) are directly supported by their own article.
On all the experiments of the baselines, we take all the source samples of different domains from trainset into account. And types of the backbone are all pretrained on ImageNet21k(Deng et al., 2009).
As for based on model-parameter-weighting few-shot methods: MCW(Lee et al., 2019) and H-ensemble(Wu et al., 2024), since they have not taken experiments on DomainNet and Office-Home datasets with ViT-Small backbone and 10-shot per class training samples, we take it by changing the backbone and samples condition while maintaining other configurations supported by their own work. And for fairness and efficiency, the well-trained source models from different domains of these methods are directly equipped with the same models trained by the first stage of our OTQMS method respectively.
As for based on samples few-shot methods: We exam the WADN(Shui et al., 2021) method under the condition of limited label on target domain as this setting is similar to ours. And we also change the backbone and samples condition while maintaining other configurations of its report to realize the experiments on DomainNet and Office-Home datasets. It is apparent that the settings of MADA(Zhang et al., 2024) are quite different of ours, leveraging all the labeled source data and all the unlabeled target data with the few-shot labeled ones that are difficult to classify, which means all the target samples have been learned to some extent. So we not only decrease the labeled target samples to 10-shot per class but also the unlabeled target samples. And since MADA is the most SOTA and comparable method, we realize it with our fair 10-shot setting under both ViT-S and ResNet50 backbone on DomainNet and Office-Home datasets while maintaining other configurations of its report.
C.5 Data efficient test details in 5.5 .
@@ C I P Q R S Avg C I P Q R S Avg
MADA(Zhang et al., 2024) ViT-S 8.70 13.39 8.31 4.02 19.76 21.32 12.58 6.40 12.52 6.29 3.22 13.19 18.18 9.97
MADA(Zhang et al., 2024) ResNet50 7.83 8.21 4.57 3.69 4.35 8.78 6.24 7.56 9.58 4.56 4.05 4.21 9.58 6.59
@@
MADA(Zhang et al., 2024) ViT-S 15.98 16.41 15.93 15.21 16.80 16.87 16.35 11.07 11.74 11.05 10.38 11.79 12.11 11.51
MADA(Zhang et al., 2024) ResNet50 15.87 15.92 15.33 15.12 15.29 15.99 15.65 11.23 11.47 10.73 10.61 10.65 11.47 11.10
@@
C.6 Domain choosing analysis details.
@@
Appendix D Method to get $s^{*}$ and $\underline{\alpha}^{*}$ which minimize (14) in Theorem 4.6
The minimization problem of K-L measure (14) is
@@
Due to the complex constraints between $s$ and $\alpha$, obtaining an analytical solution to this problem is challenging. Therefore, we propose a numerical approach to get the optimal solution. Specifically, we perform a grid search over the feasible domain of $s$. For each candidate $s^{\prime}$, we compute the optimal $\alpha^{\prime}$ under the constraint $A(s^{\prime})$, which is a $K\times K$ quadratic programming problem with respect to $\alpha$. The K-L measure (14) value is then evaluated using $(s^{\prime},\alpha^{\prime})$. After completing the search over $s$, the optimal values $s^{*}$ and $\alpha^{*}$ correspond to the pair $(s^{\prime},\alpha^{\prime})$ that minimizes the K-L measure.

### v3

@@ C.3 Proof of Proposition 5
Combining with Lemma  1, we will know that the K-L measure is
@@
C.4 Proof of Proposition 6
Similar to (C.2), we can get
[…]
@@ C.4 Proof of Proposition 6
$\sqrt{N_{0}+n_{1}}\left(\hat{{\underline{\theta}}}-E_{\hat{{\underline{\theta}}}}\right)\xrightarrow{d}\mathcal{N}\left(0,J({\underline{\theta}}_{0})^{-1}\right)$ (62)
So we can know that $\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-E_{\hat{{\underline{\theta}}}}\right)^{2}\right]$ has the limit
@@
Combining (C.4), (63) and (64), we can know that the K-L measure is
@@ C.5 Proof of Theorem 7
$\sqrt{N_{0}+\sum\limits_{i=1}^{K}n_{i}}\left(\hat{{\underline{\theta}}}-E_{\hat{{\underline{\theta}}}}\right)\xrightarrow{d}\mathcal{N}\left(0,J({\underline{\theta}}_{0})^{-1}\right)$ (67)
So we can know that $\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-E_{\hat{{\underline{\theta}}}}\right)^{2}\right]$ has the limit
@@
$E_{\hat{{\underline{\theta}}}}=\frac{N_{0}{\underline{\theta}}_{0}+\sum\limits_{i=1}^{k}n_{i}{\underline{\theta}}_{i}}{N_{0}+\sum\limits_{i=1}^{K}n_{i}}+O\left(\frac{1}{N_{0}+\sum\limits_{i=1}^{K}n_{i}}\right)=\frac{N_{0}{\underline{\theta}}_{0}+\sum\limits_{i=1}^{k}n_{i}{\underline{\theta}}_{i}}{N_{0}+s}+O\left(\frac{1}{N_{0}+s}\right)$ (69)
Combining (C.4), (68) and(69) , for the d-demension parameter, we can know that the K-L measure is
@@
Appendix D Experiment Details
D.1 Details information of LoRA framework experiments.
@@ Appendix D Experiment Details
D.2 Experimental Design and Model Adaptation
@@
D.3 Single-Source transfer performance details.
@@
D.4 Baselines experiments settings in Table 2.
Since the significant difference from unsupervised methods and supervised methods, the results of MSFDA[21], DATE[9] and M3SDA[19] are directly supported by their own article.
On all the experiments of the baselines, we take all the source samples of different domains from trainset into account. And types of the backbone are all pretrained on ImageNet21k[5].
As for based on model-parameter-weighting few-shot methods: MCW[12] and H-ensemble[31], since they have not taken experiments on DomainNet and Office-Home datasets with ViT-Small backbone and 10-shot per class training samples, we take it by changing the backbone and samples condition while maintaining other configurations supported by their own work. And for fairness and efficiency, the well-trained source models from different domains of these methods are directly equipped with the same models trained by the first stage of our OTQMS method respectively.
As for based on samples few-shot methods: We exam the WADN[22] method under the condition of limited label on target domain as this setting is similar to ours. And we also change the backbone and samples condition while maintaining other configurations of its report to realize the experiments on DomainNet and Office-Home datasets. It is apparent that the settings of MADA[32] are quite different of ours, leveraging all the labeled source data and all the unlabeled target data with the few-shot labeled ones that are difficult to classify, which means all the target samples have been learned to some extent. So we not only decrease the labeled target samples to 10-shot per class but also the unlabeled target samples. And since MADA is the most SOTA and comparable method, we realize it with our fair 10-shot setting under both ViT-S and ResNet50 backbone on DomainNet and Office-Home datasets while maintaining other configurations of its report.
D.5 Data efficient test details in 5.5 .
@@ C I P Q R S Avg C I P Q R S Avg
MADA[32] ViT-S 8.70 13.39 8.31 4.02 19.76 21.32 12.58 6.40 12.52 6.29 3.22 13.19 18.18 9.97
MADA[32] ResNet50 7.83 8.21 4.57 3.69 4.35 8.78 6.24 7.56 9.58 4.56 4.05 4.21 9.58 6.59
@@
MADA[32] ViT-S 15.98 16.41 15.93 15.21 16.80 16.87 16.35 11.07 11.74 11.05 10.38 11.79 12.11 11.51
MADA[32] ResNet50 15.87 15.92 15.33 15.12 15.29 15.99 15.65 11.23 11.47 10.73 10.61 10.65 11.47 11.10
@@
D.6 Domain choosing analysis details.
@@
Appendix E Method to get $s^{*}$ and $\underline{\alpha}^{*}$ which minimize (14) in Theorem 7
The minimization problem of proposed measure (14) is
@@ Appendix E Method to get $s^{*}$ and $\underline{\alpha}^{*}$ which minimize (14) in Theorem 7
Due to the complex constraints between $s$ and $\alpha$, obtaining an analytical solution to this problem is challenging. Therefore, we propose a numerical approach to get the optimal solution. Specifically, we perform a uniform-step exhaustive search over the feasible domain of $s$. For each candidate $s^{\prime}$, we compute the optimal $\alpha^{\prime}$ under the constraint $A(s^{\prime})$, which is a $K\times K$ quadratic programming problem with respect to $\alpha$. The proposed measure (14) value is then evaluated using $(s^{\prime},\alpha^{\prime})$. After completing the search over $s$, the optimal values $s^{*}$ and $\alpha^{*}$ correspond to the pair $(s^{\prime},\alpha^{\prime})$ that minimizes the proposed measure.
Appendix F Analysis on Domain-specific Transfer Quantity
To understand the domain preference of OTQMS, we visualize the proportion of each source in the selected samples for each target domain. As shown in Figure 5(b), when the target domain is Clipart, OTQMS primarily leverages samples from Real, Painting, and Sketch. In addition, Quickdraw contributes minimally to any target domain. These observations align with the findings in [19]. To further clarify the selection process of OTQMS during training, we visualize it in Figure 5(a). We observe that in the Office-Home dataset, Clipart initially selects all source domains but later focuses on Art and Real World. In the DomainNet dataset, Sketch predominantly selects Clipart, Painting, and Real throughout the training process.
Refer to captionRefer to caption
(a) Domain Selection
(b) Domain preference